OpenStudy (anonymous):

Simplify log(10^3*x^7)

OpenStudy (campbell_st):

start by using the log law for multiplication... \[\log_{a}(ac) = \log_{a}(b) + \log_{a}(c)\] next you need to log law for powers \[\log_{a}(b^c) = c \times \log_{a}(b)\] then if you are using base 10 logs use the law \[\log_{10}(10^a) = a \times \log_{10}(10) = a\] hope this helps

OpenStudy (anonymous):

So, is it 3x^7?

OpenStudy (campbell_st):

not quite its 3 + 7log(x)

OpenStudy (anonymous):

Do you think you could walk me through it again? Im still not really understanding it

OpenStudy (campbell_st):

ok... multiplication law \[\log(10^3) + \log(x^7)\] power law \[3 \log(10) + 7 \log(x)\] lastly same base \[3 + 7 \log(x)\]

OpenStudy (anonymous):

Oh dude, thank you. I got it now. You're awesome.